# Remote Rowhammer Attack using Adversarial Observations on Federated Learning Clients

### v1

@@ Abstract
Federated Learning (FL) has the potential for simultaneous global learning amongst a large number of parallel agents, enabling emerging AI such as LLMs to be trained across demographically diverse data. Central to this being efficient is the ability for FL to perform sparse gradient updates and remote direct memory access at the central server. Most of the research in FL security focuses on protecting data privacy at the edge client or in the communication channels between the client and server. Client-facing attacks on the server are less well investigated as the assumption is that a large collective of clients offer resilience.
Here, we show that by attacking certain clients that lead to a high frequency repetitive memory update in the server, we can remote initiate a rowhammer attack on the server memory. For the first time, we do not need backdoor access to the server, and a reinforcement learning (RL) attacker can learn how to maximize server repetitive memory updates by manipulating the client’s sensor observation. The consequence of the remote rowhammer attack is that we are able to achieve bit flips, which can corrupt the server memory. We demonstrate the feasibility of our attack using a large-scale FL automatic speech recognition (ASR) systems with sparse updates, our adversarial attacking agent can achieve around 70% repeated update rate (RUR) in the targeted server model, effectively inducing bit flips on server DRAM. The security implications are that can cause disruptions to learning or may inadvertently cause elevated privilege. This paves the way for further research on practical mitigation strategies in FL and hardware design.
Index Terms:
Federated Learning, Rowhammer Attack, Fault Injection, Reinforcement Learning
@@ I-A Motivation
Recent advances in artificial intelligence (AI), particularly in federated learning (FL), have significantly enhanced privacy preservation by enabling distributed model training across numerous client devices without sharing raw data [1]. However, the increasing complexity and scale of these distributed AI systems also expose them to novel and sophisticated threat vectors that transcend traditional adversarial machine learning techniques [2]. Among hardware security attacks, the Rowhammer attack stands out due to its potency in inducing bit-flips within Dynamic Random Access Memory (DRAM) through targeted electromagnetic interference from frequent activations of adjacent rows [3]. Despite various industry-standard countermeasures, such as Target Row Refresh (TRR), recent studies highlight persistent risks, demonstrating that modern DRAM modules remain vulnerable to advanced Rowhammer attacks [4, 5].
Figure 1: Framework of our proposed threat vector. a) a PPO agent generates b) adversarial waveforms that interfere the inputs of c) client sensors of FL, resulting the clients to send d) clustered updates, which could trigger Rowhammer attack on e) server DRAM
Motivated by the following complementary observations, we explore the potential of combining physical adversarial techniques with reinforcement learning (RL) methods to orchestrate Rowhammer bit-flips remotely in FL environments. On the one hand, existing Rowhammer attack strategies typically presume direct or indirect software-level access to victim hardware [6, 7, 8]. This assumption limits their practical applicability, particularly in scenarios where adversaries have no explicit system-level privileges or direct remote memory access. While on the other, physical adversarial attacks have demonstrated the feasibility of using carefully crafted electromagnetic or acoustic perturbations to deceive AI-driven sensor inputs without requiring direct digital intrusion [9, 10].
Specifically, modern FL systems deployed at scale frequently integrate efficiency optimizations, such as sparse gradient updates, pinned-memory regions, huge pages, and Remote Direct Memory Access (RDMA), which significantly increase their vulnerability to carefully crafted physical perturbations. We hypothesize that these optimizations, while beneficial for performance, inadvertently facilitate indirect Rowhammer attacks by enabling precisely controlled memory access patterns favorable to inducing DRAM bit-flips. Consequently, an adversary could exploit federated clients by interfering their sensor inputs with maliciously crafted noises (e.g., acoustic or electromagnetic signals) to indirectly manipulate server-side memory access patterns, ultimately causing disruption and degradation of FL system performance without direct access to victim hardware.
@@ I-B Contributions
In this paper, we propose a novel threat vector that leverages a reinforcement learning-based Proximal Policy Optimization (PPO) agent to induce Rowhammer bit-flips on federated learning servers indirectly via physical interference at client devices. The PPO agent generates stealthy adversarial waveforms projected into client sensors (e.g., microphones), triggering clustered model parameter updates on the server. These updates, amplified through efficiency optimizations such as RDMA, pinned memory, and sparse gradient updates, achieve the activation frequency and regularity necessary to induce Rowhammer attacks on server-side DRAM.
@@
    Novel Attack Vector: We introduce the first physical-domain-driven Rowhammer attack vector specifically targets federated learning (FL) systems. Unlike traditional Rowhammer attacks, our approach does not require direct access or explicit control over victim hardware or software resources, greatly expanding its potential attack surface.
@@
    PPO-based Attack Generation Framework: We design a two-stage reinforcement learning framework that generates stealthy adversarial waveforms, ensuring sustained and clustered parameter updates within FL models, which consequently induce Rowhammer bit-flips.
@@
    Vulnerability Analysis: We perform experimental evaluations using realistic federated learning scenarios with popular automatic speech recognition (ASR) models and datasets. Our experiments demonstrate that Rowhammer bit-flips can be induced on FL systems enhanced with common efficiency optimizations through adversarial interference on client sensors.
@@ I-C Organization
The remainder of this paper is organized as follows. Section II provides a relevant background on DRAM architecture, Rowhammer vulnerabilities, FL security, and physical adversarial attacks. Section III formally describes the proposed attack vector, including motivation, feasibility analysis, and attack formulation. Section IV details the experimental setups, datasets, and evaluation methodology. Section V presents extensive experimental results, analyzing the effectiveness of the PPO-based adversarial attack framework. Section VI discusses potential mitigation techniques and recommendations to counteract the proposed threat. Finally, Section VII concludes the paper and outlines directions for future research.
@@ II-A DRAM Architecture and Rowhammer Attack
Figure 2: Hierarchical organization of a DRAM system, depicting the structural composition from the memory controller and channel architecture through modules, ranks, and chips, down to individual bank-level components (row decoder, sense amplifier/row buffer) and the underlying DRAM cell structure (wordline, access transistor, capacitor). The schematic illustrates the multi-tiered interconnectivity and functional partitioning critical to DRAM operation, spanning system-level interfaces to transistor-level charge storage mechanisms.
@@
[…]
@@ II-B2 Efficiency Optimization
Federated Learning (FL) systems, especially those deployed in high-performance computing (HPC) facilities and data centers, often incorporate efficiency-oriented techniques such as sparse gradient updates, page-locked memory, and Remote Direct Memory Access (RDMA) for accelerated communication and reduced overhead.
  • Sparse updates in federated learning reduce communication overhead by transmitting only a subset of model parameter changes that exceed a chosen threshold or rank among the largest by magnitude. Clients compute local gradients and select the most significant updates to send to the server which aggregates these sparse contributions to refine the global model. This approach can significantly lower bandwidth requirements and accelerates training rounds, especially for federations of large number of clients.
  • Page-lock memory, also referred to as pinned-memory, marks a region of memory as ’pinned’, keeping the section of memory resident in RAM at all times, hence improves the efficiency of program by avoiding delays caused by relocation or page-out of memory. In modern GPU based deep learning scenario, pinned-memory locks a section of DRAM to enable asynchronous transfers between DRAM and VRAM on GPUs, which significantly improves efficiency in data loading and training throughput.
  • Huge Pages are a memory management technique that allows the operating system to use larger page sizes than the default 4KB, reducing the number of page table entries, which can improve translation lookaside buffer (TLB) hit rates and reduce the overhead of memory management. This can lead to performance improvements in applications that have large memory footprints and high memory access rates, such as high-performance computing (HPC) applications. For AI workloads, variables such as model parameters, data are often too large for default page size, hence huge pages can significantly improve the performance of AI workloads.
  • DMA and RDMA Direct Memory Access (DMA) is a feature of computer systems that allows certain hardware subsystems to access main system memory (RAM) independently of the central processing unit (CPU). Remote Direct Memory Access (RDMA) is a specialized form of DMA that enables high-speed data transfers between networked computers by allowing data to be transferred directly from the memory of one computer to the memory of another without involving the CPU. RDMA is commonly used in high-performance computing (HPC) and data center environments to accelerate data transfers and reduce latency. By bypassing the CPU and operating system, RDMA can achieve latencies as low as 1-$2\,\mu\text{s}$ and bandwidths exceeding 200 Gbps, making it ideal for applications that require high-speed, low-latency data transfers, such as machine learning training and distributed computing. In FL systems, RDMA can significantly reduce communication overhead and improve scalability by enabling efficient aggregation and distribution of model parameters between central servers and distributed clients. By minimizing synchronization delays and network bottlenecks, RDMA can enhance the overall training performance and accelerate convergence in FL systems, particularly in large-scale deployments involving hundreds or thousands of clients.
@@ II-C Physical Adversarial Attacks
Deep learning has revolutionized a wide range of applications, extending from computer vision to audio processing, enabling technologies such as autonomous driving, surveillance, biometric authentication, robotics, and voice-controlled systems. However, the reliability and robustness of these systems have been called into question due to adversarial attacks. These attacks involve purposefully crafted perturbations, which are often imperceptible to human observers, that can mislead or corrupt a model’s predictions. Originally explored in the digital domain, adversarial examples have since transitioned into the physical world, maintaining their deceptive properties even after real-world transformations like printing, photographing, or acoustic playback [24, 25, 26].
In the context of computer vision, early works by Szegedy et al. [24] and Goodfellow et al. [25] unveiled the vulnerability of neural networks to subtle pixel-level perturbations. Kurakin et al. [26] and Eykholt et al. [9] extended these findings into the physical domain, revealing that adversarial examples remain effective when transformed into physical objects or printed images. Consequently, research has demonstrated that adversarial patches [27], malicious alterations to wearable items, and tampered traffic signs [28] can consistently deceive image classification and object detection models. Such attacks often operate under the radar of human vision and can be engineered to withstand environmental noise and viewpoint changes [29, 30, 31].
Adversarial attacks have also become common in the audio domain. Speech recognition and voice-activated systems are increasingly critical in applications ranging from smart home devices to security systems. Adversarial perturbations can be embedded in audio signals to cause automatic speech recognition (ASR) models to misinterpret commands or produce incorrect transcriptions. These perturbations can be made so subtle that they remain inaudible to humans. For instance, Carlini and Wagner [10] demonstrated targeted adversarial examples against speech-to-text systems, while Yuan et al. [32] and Qin et al. [33] introduced methods for embedding covert commands within naturalistic audio clips. As with visual adversaries, these audio-based attacks persist in the physical domain, maintaining their effectiveness when played through speakers and recorded by microphones, thus bypassing human scrutiny and threatening the security of voice-controlled systems.
III Attack Vector Formulation
III-A Inspiration
@@
  1. 1.
    Stability in Physical Address: In FL workflows, training data and model parameters are assigned to locked huge pages to maximize data transfer efficiency between CPU and GPU. By allocating the buffer as a huge page, fewer TLB entries are needed, reducing address-translation overhead, while page-locking guarantees that the memory cannot be paged out or moved. DMA/RDMA then enables the device to read or write directly to this pinned, contiguous memory region. When employing these three techniques simultaneously, the operating system ensures that the allocated memory region remains both physically contiguous and non-swappable, effectively ’locking in’ a stable translation to a fixed physical address, especially when considering a short period of time such as the typical refresh period of 64ms window for typical rowhammer attacks. Therefore, potential attacker can lock the target row via the FL workflows, e.g., locking a physical DRAM row via specific continuous parameters of the global model.
  2. 2.
    Minimal Cache Interference: FL systems that employ DMA or RDMA and non-temporal memory accesses to optimize distributed parameter exchange enable more frequent and predictable direct DRAM accesses. RDMA’s zero-copy semantics let attackers initiate high-frequency remote memory reads/writes that bypass CPU caches on a targeted node’s DRAM. Additionally, sparse updates—commonly used in FL to reduce bandwidth—focus access patterns on a subset of model parameters. By continuously hammering these frequently accessed parameters or their associated pinned-memory regions, attackers can ensure sustained direct DRAM access while minimizing cache interference. Non-temporal instructions, often integrated for performance reasons, further reduce cache pollution, promoting repeated DRAM row activations.
  3. 3.
    Sustained High Rate of Row Activations: With the reduced latency and jitter in memory accesses by above techniques, potential attackers can achieve high hammering frequencies through the large number of clients with sparse gradient updates. Specifically, the gather-scatter mechanism, commonly used to efficiently aggregate sparse gradient updates by collecting and redistributing only the most significant gradients (e.g., the top 0.05%), further minimizes data transfer overhead. This concentration on a small subset of parameters not only enhances communication efficiency but also maintains consistent memory access patterns, reinforcing the repeated activation of targeted memory rows required for effective Rowhammer exploitation.
TABLE I: Common ASR models with their sizes and computation precision.
Model Params(M) Individual Tensors Precision $H_{\text{max}}(0.0005)$
Conformer-CTC-S [34] 8.7 480 INT8 296K
QuartzNet 5x5 [35] 6.7 130 INT8 384K
Citrinet 256 [36] 10.3 635 INT8 250K
Squeezeformer-XS  [37] 9.0 480 INT8 286K
Eff. Conformer-CTC [38] 13.2 520 INT8 196K

### v2

@@ Abstract
Federated Learning (FL) has the potential for simultaneous global learning amongst a large number of parallel agents, enabling emerging AI systems to be trained across demographically diverse data. Central to this being efficient is the ability for FL to perform sparse gradient updates and remote direct memory access at the central server. Most of the research in FL security focuses on protecting data privacy at the edge client or in the communication channels between the client and server. Client-facing attacks on the server are less well investigated as the assumption is that a large collective of clients offer resilience. Here, for the first time, we show that by attacking clients that lead to repetitive clustered memory updates in the server, we can remote initiate a rowhammer attack to corrupt the server memory with induced bit-flips. A reinforcement learning (RL) attacker without backdoor access to the server can learn how to maximize server repetitive clustered memory updates by manipulating the clients’ sensor observation. We demonstrate the feasibility of our attack using FL systems with sparse updates, tasked for automatic speech recognition (ASR) as well CIFAR10 classification, our adversarial attacking agent can achieve around 70% repeated update rate (RUR) in the targeted server model, effectively inducing bit flips on server DRAM. The security implications are that can cause disruptions to learning or may inadvertently cause elevated privilege. This paves the way for further research on practical mitigation strategies in FL and hardware design.
@@ I-A Motivation
Figure 1: Framework of our proposed attack vector. a) a PPO agent generates b) adversarial noises in the physical domain in the form of image patterns or waveforms that interfere the inputs of c) client sensors of FL, resulting the clients to send d) repeated clustered updates, which could trigger e) Rowhammer attack (bit-flips) on server DRAM.
Recent advances in artificial intelligence (AI), particularly in federated learning (FL), have significantly enhanced privacy preservation by enabling distributed model training across numerous client devices without sharing raw data [1]. However, the increasing complexity and scale of these distributed AI systems also expose them to novel and sophisticated threat vectors that go beyond traditional adversarial machine learning techniques [2]. Among hardware security attacks, the Rowhammer attack stands out for its ability to induce bit flips in Dynamic Random Access Memory (DRAM) through electromagnetic interference caused by frequent activations of adjacent rows [3]. Despite industry standard countermeasures such as Target Row Refresh (TRR), recent studies continue to show that modern DRAM modules remain vulnerable to advanced Rowhammer attacks [4, 5].
Motivated by the following complementary observations, we explore the possibility of combining physical adversarial techniques with reinforcement learning (RL) methods to orchestrate Rowhammer bit flips remotely in FL environments. On the one hand, existing Rowhammer strategies typically assume some form of software level access to victim hardware [6, 7, 8]. This assumption substantially limits their practical relevance in FL settings, where servers are usually difficult for potential aggressors to reach, especially at the hardware level. This challenge is expected to become even more pronounced in future deployments of FL over 6G radio access networks (RAN), where the server is likely to be placed within the central unit function and protected against backdoor entry or direct attacks. On the other hand, physical adversarial attacks have shown that carefully crafted electromagnetic or acoustic perturbations can mislead AI driven sensor inputs without direct digital intrusion [9, 10].
Specifically, large scale FL systems often incorporate efficiency optimizations such as sparse gradient updates, pinned memory regions, huge pages, and Remote Direct Memory Access (RDMA), which may unintentionally increase susceptibility to carefully crafted physical perturbations. We hypothesize that these optimizations, while beneficial for performance, can facilitate indirect Rowhammer attacks by enabling finely controlled memory access patterns that are favorable for inducing DRAM bit flips. As a result, an adversary may exploit federated clients by perturbing their sensor inputs in the physical world with maliciously crafted signals, such as acoustic or electromagnetic noise, to indirectly influence server side memory access patterns and ultimately disrupt or degrade FL system performance without requiring direct access to the protected server hardware.
@@ I-B Contributions
In this paper, we propose a novel threat vector, see Fig. 1, that leverages a reinforcement learning-based Proximal Policy Optimization (PPO) agent to induce Rowhammer bit-flips on federated learning servers indirectly via physical interference at client devices. The PPO agent generates stealthy adversarial waveforms or image patches to be capured by client sensors (e.g. microphones, cameras), triggering clustered model parameter updates on the server. These updates, conveyed by efficiency optimizations such as RDMA, pinned memory, and sparse gradient updates, achieve the activation frequency and regularity necessary to induce Rowhammer attacks on server-side DRAM.
@@
    Novel Attack Vector: We introduce the first physical-domain-driven remote Rowhammer attack vector specifically targets federated learning (FL) systems. Unlike traditional Rowhammer attacks, our approach does not require direct access or explicit control over victim hardware or software resources, greatly expanding its potential attack surface.
@@
    PPO-based Attack Generation Framework: We design a reinforcement learning framework that generates stealthy adversarial waveforms, ensuring sustained and clustered parameter updates within FL models, which consequently induce Rowhammer bit-flips.
@@
    Vulnerability Analysis: We perform experimental evaluations using realistic federated learning scenarios with popular automatic speech recognition (ASR) models and computer vision models with common open datasets. Our experiments demonstrate that Rowhammer bit-flips can be induced on FL systems enhanced with common efficiency optimizations through adversarial interference on client sensors.
@@ I-C Organization
The remainder of this paper is organized as follows. Section II provides a relevant background on DRAM architecture, Rowhammer vulnerabilities, FL security, and physical adversarial attacks. Section III formally describes the proposed attack vector, including motivation, feasibility analysis, and attack formulation. Section IV details the experimental setups, datasets, and evaluation. Section V presents experimental results, analyzing the effectiveness of the PPO-based adversarial attack framework. Section VI discusses the practical implications and limitations of our approach, mitigation techniques and recommendations to FL system design, as well as outlining directions for future research. Finally, Section VII concludes the paper.
@@ II-A DRAM Architecture and Rowhammer Attack
Figure 2: Memory allocation of FL CNN server model with PyTorch (Process-I), and access pattern at communication when receiving and processing sparse updates from clients via RDMA (Process-II). I. the CNN model consists of a hierarchical python modules, taking the ’conv 1’ layer from ’conv block 4’ as example, the tensor is breakdown into metadata and values before put into untyped storage in allocated memory, which is set to page-lock backed with Transparent Huge Pages, and mapped to a physical address of a few rows. II. When the server receive a queue of sparse updates from clients through RDMA, each message is broken down and reconstructed into a COO Tensor to be processed through the aggregation accumulators and writeback buffers of each layer tensor of the server model, which in the scenario of clustered sparse updates, these buffers (marked in red) are repeatedly accessed, serving as hammer row, potentially causing bit-flips in adjacent rows
@@
Furthermore, the reason why the Rowhammer attack requires rapid and frequent access to the aggressor row lies in the DRAM refresh mechanism. DRAM performs periodic refresh operations to maintain the charge state of all memory cells. Typically, this refresh period is $64~\text{ms}$ or $128~\text{ms}$, but under high-temperature conditions, it could be reduced to $32~\text{ms}$. In each refresh period, the MC typically issues 8192 REF commands, which sequentially refresh all rows on the DRAM module. Taking a $64~\text{ms}$ refresh period as an example, the average interval between each REF command is $7.8~\mu\text{s}$. If the number of activations on the aggressor row fails to effectively cause charge leakage in the victim row before its next refresh, the attacker must accumulate more activations again in next refresh period to execute Rowhammer attack.
@@
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Ensuring Effective Activation Rate and Count: After determine all above conditions, attackers should ensure enough activations to induce bit-flips. There are two key factors to ensure the successful execution of the attack. One is maintaining an effective activation rate. Even though the standard tRC (Row Cycle Time) is only $46~\text{ns}$, the average interval between two activation instructions should be around $220~\text{ns}$, considering the influence of other instructions. The other factor is accumulating a sufficient number of activations to ensure that the aggressor rows are activated enough times to induce bit flips. Additionally, when TRR are present, the dummy rows used to bypass TRR must also be accounted for in the total activation count. The total number of activations should not exceed the maximum allowable activations within a single refresh period ($64~\text{ms}$) to avoid DRAM refresh interventions.
@@ II-B1 Security Threats in Federated Learning
These threats are traditionally studied at the machine learning level, where the goal is to compromise model utility, implant hidden behavior, or extract private information. Our perspective is slightly different. We consider whether the same attacker controlled update channel can also be used to influence low level system behavior, particularly memory access patterns on the server, thereby creating conditions that are relevant to Rowhammer style attacks.
@@ II-B2 Efficiency Optimization
Federated learning systems deployed in high performance computing facilities and data centres often adopt aggressive communication and memory optimizations to reduce training overhead. Among the most relevant are sparse updates, page locked memory, huge pages, and DMA or RDMA based data transfer. While these mechanisms are introduced for performance reasons, they can also shape memory access behavior in ways that become security relevant.
  • Sparse updates reduce communication overhead by transmitting only a subset of model parameter changes, such as those whose magnitudes exceed a threshold or rank among the largest entries. In a typical federated learning workflow, each client computes local gradients and sends only the selected updates to the server, which then aggregates these sparse contributions into the global model. This design can substantially reduce bandwidth consumption and accelerate training, especially when the federation contains a large number of clients. At the same time, repeatedly focusing communication on a small subset of parameters can make memory access patterns more concentrated and predictable, which is potentially relevant from an attack perspective.
  • Page-lock memory, also referred to as pinned memory [24], allows application-allocated virtual memory to be fixed to physical pages, preventing it from being swapped out or migrated by the operating system and thus avoiding additional memory remapping overhead. It is commonly used in DMA/RDMA operations and high-speed data transfers in GPU-based systems. In federated learning, pinned memory is typically adopted on the server side to improve the efficiency of receiving model parameters transmitted from clients. From a system security perspective, however, during the pinned period, the parameters transmitted by clients may repeatedly access a fixed physical memory region on the server, as illustrated in Fig. 2 (a) and (b). This behaviour provides a potential opportunity for the frequent activation of specific physical rows, which is a necessary condition for enabling Rowhammer attacks.
  • Huge Pages are a memory management mechanism that enables the operating system to use larger page sizes than the default 4KB (e.g., 2MB), thereby reducing the number of page table entries, lowering address translation overhead, and improving Translation Lookaside Buffer (TLB) efficiency [25]. This technique is commonly adopted in high-performance computing (HPC) and memory-intensive applications to enhance overall memory access performance. In federated learning scenarios, where model parameters and intermediate data can be large, the server may employ huge pages to improve the efficiency of parameter buffer accesses. From an attacker’s perspective, however, the use of huge pages leads to more contiguous physical memory allocation, which increases the likelihood that different memory access regions exhibit spatial locality in the underlying DRAM. This, in turn, may increase the probability that adjacent physical rows are involved in memory accesses, thereby providing a potential basis for satisfying the physical adjacency condition required by Rowhammer attacks.
  • DMA and RDMA Direct Memory Access (DMA) and Remote Direct Memory Access (RDMA) are widely adopted high-performance data transfer mechanisms in modern computing systems. DMA allows devices to access main memory directly without CPU involvement, while RDMA further enables memory-to-memory data transfer across networked nodes, significantly reducing latency and improving bandwidth utilization. RDMA has been extensively used in high-performance computing (HPC) and data centre environments to accelerate distributed machine learning training, particularly for efficient parameter synchronization and system scalability [26]. In practice, RDMA can achieve communication latencies as low as 1-$2\,\mu\text{s}$ and support bandwidths exceeding 100 Gbps and even up to 200 Gbps, making it well-suited for federated learning systems to reduce communication overhead between clients and servers and improve overall training efficiency. From a memory access perspective, DMA/RDMA alters the traditional data path that relies on CPU cache hierarchies, allowing data transfers to interact more directly with main memory. During continuous parameter uploads from clients, this mechanism enables the server-side memory buffer to experience more intensive and stable write operations. From an attacker’s perspective, the ability to bypass CPU caches provides a direct access path to DRAM, which is a necessary condition for Rowhammer attacks. Moreover, the combination of high throughput and low latency may further increase the access intensity to specific memory regions, potentially amplifying memory disturbance effects.
@@ II-C Physical Adversarial Attacks
Deep learning systems have been challenged across a wide range of tasks, including computer vision, speech processing, autonomous driving, surveillance, biometric authentication, robotics, and voice controlled services, by adversarial attacks, in which carefully crafted perturbations cause a model to produce incorrect predictions. While early studies focused on perturbations applied directly in the digital domain, later work showed that adversarial manipulations can also remain effective after physical world transformations [27, 28, 29].
In computer vision, adversarial perturbations have been demonstrated on printed images, adversarial patches, wearable objects, and modified traffic signs, showing that malicious inputs can preserve their effect even under viewpoint changes, lighting variation, and environmental noise [9, 30, 31, 32, 33]. In the audio domain, adversarial perturbations can be embedded into speech or music so that automatic speech recognition systems produce attacker chosen transcriptions while the signal remains natural or inconspicuous to human listeners [10, 34, 35]. These results show that adversarial manipulation is not limited to synthetic digital settings, but can be embedded into real world inputs that are naturally collected by machine learning systems.
This observation is relevant to federated learning because client side training data may originate from physical environments, such as cameras, microphones, wearable sensors, or mobile devices. An adversary who can influence these inputs may be able to steer local model updates without needing direct access to the global model or server side infrastructure. In this sense, physical adversarial examples provide a practical threat vehicle for shaping the behavior of FL clients through apparently legitimate data without tackling with authentication processes.
In this paper, however, our focus is not on the misclassification effect of physical adversarial examples themselves. Instead, we use them to motivate a broader point: attacker controlled inputs and updates in FL can influence not only model behavior, but also the system level processing of selected parameters. This perspective is important for the attack vector developed next, where repeated and targeted updates interact with efficiency optimizations to create conditions favorable to Rowhammer style exploitation.
III Methodology
III-A Attack Vector Formulation
III-A1 Predictable Memory Access Patterns
@@
  • Predictable and stable physical placement. Modern federated learning systems frequently use pinned huge pages to accelerate data movement between CPUs and GPUs. Huge pages reduce translation overhead, and page locking prevents the allocated region from being swapped out or migrated. DMA and RDMA then provide direct access to these pinned buffers. Together, these mechanisms create memory regions that are both physically contiguous and temporally stable, especially within the short time scale relevant to Rowhammer, such as a 64 ms DRAM refresh window. This stability can significantly reduce the uncertainty that attackers usually face when attempting to align accesses with a chosen physical DRAM row. In effect, the federated learning workflow itself may help an attacker maintain a target row at a fixed physical location by repeatedly interacting with selected model parameters.
  • Direct DRAM access with reduced cache disruption. Efficient distributed training often depends on RDMA, DMA, sparse synchronization, and non temporal accesses. These features are beneficial for throughput, but they also reduce the degree to which CPU caches obscure or absorb repeated memory references. RDMA zero copy transfers allow remote actors to trigger frequent reads and writes directly against the target node’s memory, while sparse updates repeatedly touch only a narrow set of parameters. This combination provides a practical way to generate focused and repeated accesses to the same pinned memory regions, thereby increasing the likelihood of direct DRAM row activations. Non temporal instructions strengthen this effect by further suppressing cache residency.
  • High frequency and persistent hammering opportunities. Successful Rowhammer attacks require a sufficiently high rate of repeated row activations. Federated learning optimizations can make such access patterns easier to sustain. Reduced latency, lower jitter, and concentrated communication around sparse gradients allow attackers to issue highly repetitive memory operations with less variability. This effect becomes even stronger in deployments with many clients, where repeated sparse contributions can collectively amplify hammering intensity. Gather scatter aggregation, especially when restricted to the most significant gradients such as the top 0.05%, repeatedly targets the same small parameter subset. From an attack perspective, this creates a natural mechanism for persistent hammering of memory regions associated with those parameters.
III-A2 Stealthy Attack Vehicle
In typical federated learning deployments, corrupting the authentication process or injecting malicious clients at scale is challenging. Yet physical interferences via adversarial noises are often overlooked. Rather than requiring sophisticated software exploits or the circumvention of cryptographic authentication protocols, an attacker only needs to alter the physical space with adversarial patches within a camera’s field of view, or by broadcasting imperceptible, high-frequency audio perturbations in the vicinity of a voice-controlled IoT device. When the FL client samples its environment, this physical adversarial noise translates into the client’s local parameter updates, infiltrate the authentication process and alter server memory access patterns.
III-A3 Reinforcement Learning Formulation
We initiate a feasibility validation of the attack vector with a reinforcement learning agent targeting a FL system using page-lock memory backed with Transparent Huge Page, and send sparse updates to the server via RDMA. The agent can observes the inputs and parameter updates of the clients, and generates adversarial noises that feeds to the clients’ input after emulation of physical embodiment and re-digitalization.
TABLE I: Target Models with Sizes and Computation Precision.
Model Params(M) Trainable Tensors Precision $H_{\text{max}}(0.1\%)$ $H_{\text{max}}(0.05\%)$
Conformer-CTC-S [36] 8.7 480 INT4 296K 592K
Squeezeformer-XS  [37] 9.0 480 286K 572K
QuartzNet 5x5 [38] 6.7 130 INT8 192K 384K
MobileNet V3 Small  [39] 2.9 142  [^4] K  [^8] K